OpenStudy (anonymous):

Updated Batting Averages are computed using the formula : A = ( B + K ) / G A= batting average B = times at batting multiplied by player's seasonal average K = hits in the new game G = times at bat plus times at bat in a new game Part 1 : Solve A = (B + K)/ G for K. Please show your work. Part 2 : Use your new formula & the values given in the table below to find the player's hits in a new game (K). Round your answer to the nearest whole number .

OpenStudy (anonymous):

Batter A B G K George 0.319 10.7 46 Jose 0.320 11.4 45 Dale 0.298 12.3 52 Michael 0.337 15.2 48

OpenStudy (anonymous):

you have to isolate K the same way that A is isolated. find K =

OpenStudy (anonymous):

first step is to multiply both sides by G AG = B + K K = AG - B then you have to build the table with this using the given A,G and B

OpenStudy (anonymous):

So would it be 46 * 0.319 = (10.7 + K ) / 46 * 46 ?

OpenStudy (anonymous):

K = AG - B K = (.319)(.46) - 10.7

OpenStudy (anonymous):

How did you get that ?

OpenStudy (anonymous):

part 1:A = (B+K)/G Multiply by G: AG = B+K. Subtract B: AG - B = K.

OpenStudy (anonymous):

your new formula is K = AG - B and you just sub your numbers into that formula

OpenStudy (anonymous):

part 2: a. A= (B+K)/G, for K (G) A = ((B+K)/G) (G) , Multiply both sides by G GA = B+K GA-B = B+K -B , Subtract B from both sides GA-B = K b. George ---> K= GA-B = (46)(0.319) - 10.7 = 3.974 = 4 Jose ----> K= GA-B= (45)(0.320) - 11.4 = 3 Dale ----> K= GA-B= (52)(0.298) - 12.3 = 3.196 = 3 Michael --> K= GA-B= (48)(.3337) - 15.2 = 0.976 = 1

OpenStudy (anonymous):

@Booisme ; I've already done that 3 times . This is the 4th time doing this assignment & The first 3 times it's been wrong .

OpenStudy (anonymous):

you did use the formula K = AG - B, didn't you ?

OpenStudy (anonymous):

Booisme is correct

OpenStudy (anonymous):

Yeah .. I've done it the exact same way & My teacher is like : " Not quite , Bella . "

OpenStudy (anonymous):

I do not see any other way to do it...sorry

OpenStudy (anonymous):

Maybe I did mess up .. Iunno . Thank you both . (:

OpenStudy (anonymous):

A = (B + K)/G .319 = (10.7 + K) / 46 multiply by 46 14.674 = 10.7 + K 14.674 - 10.7 = K 3.974 = K 4 = K even this way gives you the same answer

OpenStudy (anonymous):

That's what I get too ..

OpenStudy (anonymous):

and your teacher says it is wrong ?

OpenStudy (anonymous):

Yeah . :/

OpenStudy (anonymous):

I do not understand your teacher.....you are using the formula, plugging in the numbers, and rounding your answer to the nearest whole number. I do not know what to tell you. Talk to her again, maybe she can explain one to you, although I do not see what is wrong about it.

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

Alright , Thank you again though .

OpenStudy (anonymous):

YW :)